# Space to Policy: Scalable Brick Kiln Detection and Automatic Compliance Monitoring with Geospatial Data

### v1

@@ 6.2. Source Apportionment
Emissions can be estimated from various sources, but they do not translate to air pollution without accounting for meteorology and complex physicochemical reactions in the atmosphere. PM2.5 and PM10 are major pollutants linked to respiratory diseases in humans (Thangavel et al., 2022). Using guidance from an air quality expert, we estimated PM2.5 and PM10 concentrations in the Delhi Airshed by employing the Chemical Transport Model (CTM), CAMx (camx, 2019). CAMx utilizes meteorological inputs generated by the Weather Research and Forecasting (WRF) model (Skamarock et al., 2019). We ran the model over the Delhi Airshed on an 80 km × 80 grid, covering the months of March and April 2024.
@@ 8. Related Work
We discuss related work in three areas: object detection from aerial imagery, the application of satellite imagery for addressing UN Sustainable Development Goals (SDGs), and prior efforts on brick kiln detection.
@@ 8.2. Computer Vision on Satellite Imagery for UN SDGs
8.3. Brick Kiln Detection from Satellite Imagery
Various machine learning methods have been applied to detect brick kilns in South Asia. Nazir et al. (Nazir et al., 2020) used a gated neural network combining classification and detection tasks to identify kilns over a 3,300 km2 area with 1,300 kiln images. Lee et al. (Lee et al., 2021) applied deep learning in Bangladesh, identifying 3,345 kilns. A classical machine learning approach using Sentinel-2 imagery classified kilns in Delhi (Imaduddin et al., 2023). Haque et al. (Haque et al., 2022) analyzed heavy metal contamination in soils near brick kilns in Bangladesh. Boyd et al. (Boyd et al., 2018) employed YOLOv3 with axis-aligned bounding boxes (AABBs) to detect fixed chimney bull’s trench kilns (FCBKs) in the Indo-Gangetic plain. However, most previous studies relied on expensive high-resolution satellite imagery, significantly limiting scalability and reproducibility.
Our study marks the first large-scale effort to geo-locate and classify brick kilns by technology type (CFCBK, FCBK, Zigzag) across five states in the Indo-Gangetic Plain, spanning 520,000 km2. We identified kilns, validated our dataset with field surveys, and assessed the evolution of kiln technologies over time. Unlike prior works, we used freely available, moderate-resolution satellite imagery (4.77 m resolution) under an Education and Research license, ensuring scalability, affordability, and reproducibility.

### v3

@@ 6.2. Source Apportionment
Emissions can be estimated from various sources, but they do not translate to air pollution without accounting for meteorology and complex physicochemical reactions in the atmosphere. PM2.5 and PM10 are major pollutants linked to respiratory diseases in humans (Thangavel et al., 2022). Using guidance from an air quality expert, we estimated PM2.5 and PM10 concentrations in the Delhi Airshed by employing the Chemical Transport Model (CTM), CAMx (camx, 2019). CAMx utilizes meteorological inputs generated by the Weather Research and Forecasting (WRF) model (Skamarock et al., 2019a). We ran the model over the Delhi Airshed on an 80 km × 80 grid, covering the months of March and April 2024.
@@ 7. Discussion
In this section, we discuss the broader implications of our findings, focusing on socio-economic, environmental, and policy aspects. We explore how this research can aid decision-making in the brick kiln sector without disrupting the livelihoods of workers. Additionally, we examine how our data can enhance air quality modeling by improving emission inventories, the importance of our work with respect to policies on transitioning kilns to Zigzag technology as per MoEFCC guidelines, and the challenges of isolating brick kiln pollution from satellite data.
7.1. Socio-Economic Considerations
@@
7.2. Data Implications:
Access to accurate and up-to-date location data for brick kilns is crucial for predicting air pollution. Our method can be utilized to update kiln locations every month or season. This information can support the air quality modeling community by improving emission inventories, considering changes in kiln technology that affect emissions and dust resuspension from land use. Currently, estimating the daily and seasonal production of a brick kiln is extremely difficult, making it challenging to determine its emissions. However, with further analysis on a higher resolution imagery, this study can be extended to record the designed daily production capacity of each kiln and the number of days or weeks it operates in a year, leading to more accurate emission estimates. Production capacity can be estimated by automatically computing the trench width of kilns. Production cycle and operating status can be estimated by detecting the open fuel holes on top of the brick kilns. Over time, these holes keep moving to process the next batch of the bricks, which can give us an estimate of production in a certain duration. This approach not only enhances the precision of emission modeling but also provides valuable insights for policymakers to regulate industrial activities more effectively. Furthermore, the integration of such data could facilitate the development of targeted mitigation strategies to reduce air pollution from brick kilns.
7.3. Policy Implications
Understanding the extent of air pollution caused by brick kilns is important both in terms of location and time. Spatially, it is necessary to determine how far and wide brick kilns are spread within and around an urban area. Temporally, it is essential to know when these kilns are active and how they impact local air quality. Additionally, evaluating the compliance rates of operational brick kilns is crucial, not only by examining past data but also by tracking them in near real-time while considering the rapid pace of urbanization. The Ministry of Environment, Forests, and Climate Change (MoEFCC) has set a revised deadline requiring all brick kilns to switch to zigzag kiln technology by February 2026. Checking the compliance rate at the district level will help state regulators identify districts that are falling behind and take additional steps to improve compliance in these areas. Monitoring compliance over time (Fig 11) also provides important insights into how long it takes for a new technology to spread. This information is valuable not only for researchers studying how small industries adopt new technologies but also for policymakers in setting realistic timelines for achieving full adoption of a new technology requirement. Furthermore, continuous monitoring of compliance will enable policymakers to assess whether additional incentives or enforcement measures are needed to ensure sustained adherence to environmental standards. Such efforts are particularly critical in regions experiencing rapid urbanization, where the environmental and health impacts of industrial activities are magnified.
7.4. Challenges in Isolating Brick Kiln Pollution from Satellite Data
Brick kilns primarily emit CO2, CO, SO2, and NOx. Satellite products, such as Sentinel-5p (Agency, 2017), provide global coverage of air quality and greenhouse gases. Recent studies (Trenchev et al., 2023; Meier, 2023) have demonstrated the effectiveness of Sentinel-5p in identifying emissions from power plants and coal mines. However, isolating brick kiln emissions from satellite data presents significant challenges due to the following factors:
Low spatial resolution: The spatial resolution of Sentinel-5p is 7 x 7 km2, whereas a typical brick kiln measures approximately 200 x 200 m. This coarse resolution makes it difficult to detect the opening or closure of individual brick kilns using emission data. Furthermore, the large footprint of satellite measurements often includes multiple emission sources, complicating the attribution of emissions to specific brick kilns.
Low intensity emissions: Brick kiln emissions are significantly weaker compared to those from power plants. As shown in Table 12, a single brick kiln emits only 0.12 tonnes of CO per day. In contrast, a coal-fired power plant in India emits 48 tonnes of CO per day (Guttikunda and Jawahar, 2014), which is 400 times higher than the emissions from a brick kiln. The combination of low-resolution data and low-intensity emissions makes it particularly challenging to isolate brick kiln contributions from satellite observations. This limitation highlights the need for alternative or complementary approaches, such as higher-resolution satellite systems or ground-based measurements, to better quantify brick kiln pollution.
However, this does not indicate that brick kiln pollution is significantly lower than that of power plants, due to the fact that the number of brick kilns is much higher compared to the number of power plants. For example, as of 2025, Uttar Pradesh has 19 major coal-based power plants555https://en.wikipedia.org/wiki/Category:Coal-fired_power_stations_in_Uttar_Pradesh and around 20,000 brick kilns (approximately 1000x), which indicates that the contribution of brick kilns to CO generation is more than twice that of power plants in Uttar Pradesh.
@@ 8. Related Work
In this section, we discuss related work in the following areas: object detection from aerial imagery, the application of satellite imagery for addressing UN Sustainable Development Goals (SDGs), impact of brick kilns on air pollution and health, brick kiln detection and air quality modeling.
@@ 8.2. Computer Vision on Satellite Imagery for UN SDGs
8.3. Impact of Brick Kilns on Air Pollution and Health
In this subsection, we highlight studies related to the impact of brick kilns on air pollution and health. Guttikunda (2009, 2008) studied the impact of brick kiln air pollution on health in Dhaka, Bangladesh. Their research revealed that air pollution from brick kilns contributes to nearly 5000 premature deaths annually in Dhaka city. Guttikunda et al. (2013) estimated that brick kilns account for 23 to 30% of air pollution in various clusters within the Greater Dhaka region, Bangladesh. Furthermore, Guttikunda et al. (2014) demonstrated that transitioning to Zigzag and Hoffman technologies can reduce PM emissions by 40% and 90%, respectively, compared to the traditional Fixed Chimney Bull’s Trench Kiln (FCBK) technology. Le and Oanh (2010) conducted a study in Northern Vietnam, estimating emission factors of brick kilns and using the Industrial Source Complex Short-Term (ISCST3) dispersion model to assess PM, CO, and SO2 concentrations. Their findings indicated that SO2 emissions exceeded National Ambient Air Quality Standards over 63 km2 of a 100 km2 modeled region during 2006-07. Similarly, Skinder et al. (2014) monitored emissions from a cluster of kilns in Jammu & Kashmir, finding that all measured pollutants, including oxides of sulphur (SOx), oxides of nitrogen (NOx), Respirable Suspended Particulate Matter (RSPM), and Non-Respirable Suspended Particulate Matter (NRSPM), exceeded National Ambient Air Quality Standards (NAAQS) during operational phases. Rajarathnam et al. (2014) derived emission rates from various brick kiln technologies, estimating annual PM emissions of 0.94 million tonnes. In Delhi, Guttikunda and Goel (2013) estimated that brick kilns contribute 4-17% to air pollution, attributing 300 to 2700 premature deaths and 0.24-1.02 million asthma attacks annually to brick kiln-generated air pollution. Raza and Ali (2021) assessed the lung health of brick kiln workers in Kasur district, Pakistan, finding that 78% of workers exhibited abnormal lung function. A comprehensive review by Nicolaou et al. (2024) of 104 studies revealed consistent adverse health outcomes among brick kiln workers, including reduced lung function, increased respiratory symptoms, and higher rates of musculoskeletal complaints compared to unexposed populations. These findings underscore the critical need for stricter regulations and cleaner technologies to mitigate the significant health and environmental impacts of brick kiln emissions.
8.4. Brick Kiln Detection
[…]
| Paper ID | Satellite Imagery | Classification | Area (km2) | Annotated Kilns | Detected + Verified Kilns |
| Boyd et al. (2018) | Google Earth | No | 32,000 | 1142 | 0 |
| Nazir et al. (2019) | Maxar | No | - | 102 | 0 |
| Nazir et al. (2020) | Maxar | No | 3,300 | 521 | 3392 |
| Misra et al. (2020) | Sentinel-2 | No | 6,400 | 282 | 1564 |
| Lee et al. (2021) | Maxar | FCBK, Zigzag | 148,460 | - | 6978 |
| Boyd et al. (2021) | Airbus | FCBK | 1,551,997 | 1142 | 66455 |
| Paul et al. (2022) | Maxar | FCBK | - | 641 | - |
| Hamdani et al. (2024) | Sentinel-2, Google Maps | FCBK, Zigzag | 518,000 | 670 | 11000 |
| Mondal et al. (2024) | Google Maps | No | 276,000 | 1042 | 7477 |
[…]
8.5. Air Quality Modeling
Brick kiln detection helps update the emission inventory, which is a critical component used in air quality modeling. In this section, we highlight notable research work in air quality modeling. Agarwal et al. (Agarwal et al., 2024) utilized NOAA’s HYSPLIT model and NASA’s GEOS-CF dataset to track particle movements and identify air pollution sources in Delhi. Our work can enhance the accuracy of the emission inventory used in the GEOS-CF modeling framework, thereby improving the precision of Agarwal et al. (2024). Guttikunda et al. (2019a) conducted gridded chemical transport modeling of air pollution using the CAMx model (camx, 2019), incorporating emission inventory and meteorology data modeled with WRF (Skamarock et al., 2019b). Similar studies have employed models such as ISCST3 (Le and Oanh, 2010), AERMOD (Kanabkaew and Buasing, 2015), and CALPUFF (Xie et al., 2021b). Another research direction (Cheng et al., 2018; Patel et al., 2022; Hu et al., 2023; Liang et al., 2023; Wang et al., 2025) focuses on modeling air quality data using ground-based sensor observations and additional predictors, including meteorological data, Points of Interest (PoIs), and road networks.
@@ 9. Limitations and Future Work
We explored the out-of-region performance of models in Section 4.4 and found that models perform poorly in out-of-region settings. This limitation underscores the need for more robust generalization capabilities in current architectures. Future work can leverage domain adaptation techniques to address this challenge, potentially reducing exclusion errors and improving model applicability across diverse scenarios. Additionally, exploring adaptive methods that account for regional differences could enhance the reliability of models in real-world applications.